### v1

# Investigation of $T_{cs0}^{*}(2870)^{0}$ in $pp$ collisions at $\sqrt{s}$ = 7 TeV with the PACIAE model

Abstract
We have used the parton and hadron cascade model PACIAE together with the Dynamically Constrained Phase-space Coalescence model (DCPC) to study the $T_{cs0}^{*}(2870)^{0}$ production in $pp$ collision at $\sqrt{s}$ = 7 TeV, in line with the LHCb observation of $T_{cs0}^{*}(2870)^{0}$ in the $B^{-}\to D^{-}D^{0}K^{0}_{S}$ decays in $pp$ collisions at $\sqrt{s}$ = 7, 8, and 13 TeV [PRL 134(2025)101901]. The final hadronic states of the $pp$ collisions at $\sqrt{s}$ = 7 TeV are first simulated by the PACIAE model. Three sets of $T_{cs0}^{*}(2870)^{0}$ candidates are then recombined by the DCPC model using the constituent meson pair of $D^{0}K^{0}_{S}$, $D^{+}K^{-}$, and $D^{-}K^{+}$ based on the above simulated final hadronic states, respectively. We calculate their rapidity distributions, transverse momentum spectra, and angular distribution between the two component mesons, as well as angular distribution between $D$ component meson and $T_{cs0}^{*}(2870)^{0}$. Our results show that the yields of three $T_{cs0}^{*}(2870)^{0}$ candidates follow the magnitude order of $D^{+}K^{-}$ $>$ $D^{-}K^{+}$ $\sim$ $D^{0}K^{0}_{S}$ (the latter two are nearly equal indeed). Similar ordering behavior is also observed in the aforementioned distributions.
journal: Physics Letters B
[…]
[first]organization=Key Laboratory of Quark and Lepton Physics (MOE) and Institute of Particle Physics, Central China Normal University,city=Wuhan, postcode=430079, country=China
[…]
[second]organization=School of Mathematics and Statistics, Wuhan Textile University,city=Wuhan, postcode=430200, country=China
[…]
[third]organization=School of Physics and Information Technology, Shaanxi Normal University,city=Xi’an, postcode=710119, country=China
[…]
[fourth]organization=China Institute of Atomic Energy, P. O. Box 275 (10),city=Beijing, postcode=102413, country=China
1 Introduction
The existence of exotic states containing more than three quarks is allowed by the quark model and quantum chromodynamics (QCD) theory [1, 2, 3]. The first exotic hadron candidate X(3872) (or $\chi_{c1}(3872)$) was experimentally observed by the Belle Collaboration in 2003 [4]. Since then more and more exotic states have been confirmed experimentally [5, 6], including $T_{cc}(3875)^{+}$ [7, 8], $T_{cc\bar{c}\bar{c}}(6900)^{0}$ [9, 10, 11], etc. In 2020, the LHCb Collaboration reported two new exotic states in the $D^{+}K^{-}$ final state of the $B^{-}\to D^{-}D^{+}K^{-}$ decays: $T_{cs0}^{*}(2870)^{0}$ and $T_{cs1}^{*}(2900)^{0}$, with quark content of $cs\bar{u}\bar{d}$ and with quantum number of $J^{P}=0^{+}$ and $1^{-}$, respectively [12, 13]. They are the first two manifestly observed open-charm tetraquark exotic states containing one single charm quark, revealing a new horizon for understanding the mechanism of quark confinement. In 2025, the LHCb collaboration confirmed the existence of $T_{cs0}^{*}(2870)^{0}$ in the $D^{0}K^{0}_{S}$ invariant mass spectrum from the new decay mode of $B^{-}\to D^{-}D^{0}K^{0}_{S}$ in $pp$ collisions at $\sqrt{s}$ = 7, 8, and 13 TeV. Unfortunately, there is no evidence for $T_{cs1}^{*}(2900)^{0}$ in the mass spectrum of $D^{0}K^{0}_{S}$ in this channel [14]. These experimental observations will deepen our understanding of the nature of $T_{cs0}^{*}(2870)^{0}$.
In analogous to other exotic states, the physical structures of $T_{cs0}^{*}(2870)^{0}$ and $T_{cs1}^{*}(2900)^{0}$ are controversial, and their nature is still unclear. In Refs. [15, 16, 17, 18, 19, 20, 21, 22] they have been suggested as compact tetraquark states. But they were interpreted as hadronic molecular states in Refs. [14, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32]. In line with Ref. [14], we assume $T_{cs0}^{*}(2870)^{0}$ is a hadronic molecular state and employ the parton and hadron cascade model PACIAE [33] together with the Dynamically Constrained Phase-space Coalescence model (DCPC [34]; PACIAE + DCPC) to study its nature and properties.
The PACIAE + DCPC model has been successfully used to describe the production of light nuclei (such as deuterons and hypertritons) in heavy-ion collisions at various center-of-mass energies [34, 35]. It has also been applied to describe the production of exotic states, e.g., X(3872) [36, 37, 38], $P_{c}$ states [39], $Z_{c}^{\pm}$(3900) [40], X(2370) [41], and G(3900) [42] in nuclear collisions.
Figure 1: A schematic diagram of PACIAE simulation for high-energy proton-proton collisions.
In this letter, the PACIAE model is first employed to simulate the final hadronic states (FHS) in $pp$ collisions at $\sqrt{s}$ = 7 TeV. The $T_{cs0}^{*}(2870)^{0}$ molecular state is then coalesced in this FHS with the component meson pair of $D^{0}K^{0}_{S}$ and $D^{+}K^{-}$ (with conjugate $D^{-}K^{+}$), by the DCPC model. These two component meson pairs correspond to the $T_{cs0}^{*}(2870)^{0}$ candidates observed by LHCb in decay channels of $B^{-}\to D^{-}D^{0}K^{0}_{S}$ and $B^{-}\to D^{-}D^{+}K^{-}$, respectively. The three sets of $T_{cs0}^{*}(2870)^{0}$ candidates, composed, respectively, of $D^{0}K^{0}_{S}$, $D^{+}K^{-}$, and $D^{-}K^{+}$ have the same quark content of $cs\bar{u}\bar{d}$. We calculate their rapidity ($y$) distributions, transverse momentum ($p_{T}$) spectra, and angular distribution between two component mesons of $D$ and $K$ ($D^{0}$ and $K^{0}_{S}$, $D^{+}$ and $K^{-}$, $D^{-}$ and $K^{+}$), as well as the angular distribution between $D$ ($D^{0}$, $D^{+}$, or $D^{-}$) component meson and $T_{cs0}^{*}(2870)^{0}$. Our results show that the yield of $T_{cs0}^{*}(2870)^{0}$ composed of $D^{+}K^{-}$ is larger the ones composed of $D^{-}K^{+}$ and $D^{0}K^{0}_{S}$. The latter two are nearly equal indeed. Similar ordering behavior is also found in all the aforementioned distributions.
2 Methodology
2.1 PACIAE model generating final hadronic state
The parton and hadron cascade model PACIAE [43, 44] based on PYTHIA [45], can be used to describe high-energy elementary particle collisions as well as nucleus-nucleus collisions. It contains four stages of the parton initialization, the parton rescattering, the hadronization, and the hadron rescattering, as shown in Fig. 1 for a $pp$ collision. The latest version of PACIAE 4.0 [33] is based on PYTHIA 8.3 [46]. In this letter, we utilize it to generate the final hadronic state in $pp$ collisions at $\sqrt{s}$=7 TeV.
In the parton initialization stage, each nucleon-nucleon (hadron-hadron) collision is executed by PYTHIA model [46] with presetting the string fragmentation turn-off temporarily. The obtained parton initial state is produced by QCD hard scattering and associated initial and final QCD radiation. Then the gluons split and the energetic quarks (antiquarks) deexcitation is implemented. In the next stage of parton rescattering, the leading-order perturbative quantum chromodynamics (LO-pQCD) parton-parton scattering cross section [47] is employed to supply the final partonic state (FPS) composed of abundant quarks and antiquarks with their four-coordinate and four-momentum. The hadronization process is then achieved via the Lund string fragmentation mechanism and/or the phenomenological coalescence model [33], and the latter is chosen in this letter. Eventually, the final hadronic state (FHS) is produced after the hadron rescattering. This FHS is composed of numerous hadrons with their four-coordinate and four-momentum.
2.2 DCPC model recombining $T_{cs0}^{*}(2870)^{0}$ candidates
The DCPC model is based on the quantum statistical mechanics [48, 49], where the yield of $N$-particle cluster is defined as
$Y_{1}=\int...\int_{E_{a}\leq H\leq E_{b}}\frac{\mathrm{d}\vec{q}_{1}\mathrm{d}\vec{p}_{1}...\mathrm{d}\vec{q}_{N}\mathrm{d}\vec{p}_{N}}{h^{3N}},$ (1)
where $E_{a}$ and $E_{b}$ denote the lower and upper energy thresholds of the cluster. The $\vec{q}_{i}$ and $\vec{p}_{i}$ are, respectively, the coordinate and momentum of the $i$-th particle. In order for the cluster to exist naturally, some dynamical constraints are introduced in the DCPC model. Thus the yield of $T_{cs0}^{*}(2870)^{0}$ composed of $D$ and $K$ component meson pair, for instance, is calculated by
$Y_{T_{cs0}^{*}(2870)^{0}}=\int...\int\delta_{12}\frac{\mathrm{d}\vec{q}_{1}\mathrm{d}\vec{p}_{1}\mathrm{d}\vec{q}_{2}\mathrm{d}\vec{p}_{2}}{h^{6}},$ (2)
where
$\delta_{12}=\begin{cases}1\quad\rm{if}\quad 1\equiv D,2\equiv K,\\ \qquad m_{0}-\Delta m\leq m_{\rm{inv}}\leq m_{0}+\Delta m,\\ \qquad R_{i}\leq R_{0},(i=1,2)\\ 0\quad\rm{otherwise}.\end{cases}$ (3)
In Eq.(3), $m_{0}$ ($\Delta m$) denotes the mass (mass uncertainty parameter) of the cluster ($T_{cs0}^{*}(2870)^{0}$ candidate). $R_{0}$ and $R_{i}$ refer to the radius of the cluster (a free parameter) and the relative distance between the $i$-th component meson and the center of mass of the cluster, respectively. The invariant mass $m_{inv}$ is defined as
$m_{\rm{inv}}=\sqrt{(E_{1}+E_{2})^{2}-(\vec{p}_{1}+\vec{p}_{2})^{2}},$ (4)
where $E_{1}$ ($E_{2}$) and $\vec{p}_{1}$ ($\vec{p}_{2}$) denote, respectively, the energy and momentum of $D$ ($K$) component meson.
2.3 Calculation routine in DCPC
We first construct a $T_{cs0}^{*}(2870)^{0}$ component meson list (consisted of $K^{+}$, $K^{-}$, $K^{0}_{s}$, $D^{+}$, $D^{-}$, and $D^{0}$) based on the hadron list in the FHS generated by the PACIAE model. According to the Eq. (3) we then construct a two-layer circulations (loops) over the component meson list to filter out any one of possible candidate among the studied three sets of $T_{cs0}^{*}(2870)^{0}$ candidates (composed of $D^{0}K^{0}_{S}$, $D^{+}K^{-}$, and $D^{-}K^{+}$). Here the parameters in Eq. (3) are set as follows: $m_{0}=2866$ MeV/$c^{2}$ (the experimentally observed mass peak of $T_{cs0}^{*}(2870)^{0}$ molecular state); $\Delta m$ equals to 28.5 MeV/$c^{2}$ (the half decay width of the $T_{cs0}^{*}(2870)^{0}$ [50]); and the $R_{0}$ is assumed to be in the range ($R_{0}$ range) of 1.0 fm $<R_{0}<$ 10.0 fm  [51, 52], or 1.0 fm $<R_{0}<$ 5.0 fm, or 1.0 fm $<R_{0}<$ 2.0 fm.
3 Results and discussion
The PACIAE 4.0 model (with coalescence hadronization) is used to generate 300 million final hadronic states for the $pp$ collisions at $\sqrt{s}$ = 7 TeV. Here we adjust only two main parameters of adj1(16) (the allowed number of deexcitation generations of the energetic quark (antiquark)) and adj1(17) (the threshold energy of deexcitation). Other parameters are default. The adj1(16) and adj1(17) are fixed by fitting the ALICE measured $\pi^{+}$, $K^{+}$, $D^{0}$ and $D^{+}$ yields at mid-rapidity region in $pp$ collisions at $\sqrt{s}$ = 7 TeV [53, 54]. The simulated yields (with fitted parameters of adj1(16)=1 and adj1(17)=2.15 GeV) and the corresponding experimental data are shown in Table 1.
Then we use above generated final hadronic states to recompose three sets of $T_{cs0}^{*}(2870)^{0}$ hadronic molecular state candidates by the DCPC model using the constituent meson pairs of $D^{0}K^{0}_{S}$, $D^{+}K^{-}$, and $D^{-}K^{+}$, respectively.
Table 1: The yields of $\pi^{+}$, $K^{+}$, $D^{0}$, and $D^{+}$ simulated by PACIAE model, compared with the ALICE data at mid-rapidity region in $pp$ collisions at $\sqrt{s}$=7 TeV [53, 54].
Particle ALICE PACIAE
$\pi^{+}$ $2.26\pm 0.10$ 2.27
$K^{+}$ $0.286\pm 0.016$ 0.284
$D^{0}$ $(8.04\pm 1.57)\times 10^{-3}$ $8.25\times 10^{-3}$
$D^{+}$ $(2.93\pm 0.72)\times 10^{-3}$ $3.06\times 10^{-3}$
The simulated $y$ (left panel) and $p_{T}$ (right panel) distributions of three $T_{cs0}^{*}(2870)^{0}$ candidates (recombined by $D^{0}K^{0}_{S}$, $D^{+}K^{-}$, and $D^{-}K^{+}$) in $pp$ collisions at $\sqrt{s}$ = 7 TeV are given in Fig. 2. In this figure, the black, blue, and red markers are the results calculated in the $R_{0}$ range of 1.0 $<R_{0}<$ 10.0 fm, 1.0 $<R_{0}<$ 5.0 fm, and 1.0 $<R_{0}<$ 2.0 fm, respectively. The triangle-up, the circle, and the triangle-down refer to, respectively, the results of the $D^{+}K^{-}$, the $D^{-}K^{+}$, and the $D^{0}K^{0}_{s}$ component meson pairs. It is observed that the yield of $T_{cs0}^{*}(2870)^{0}$ composed of $D^{+}K^{-}$ is larger than the ones composed of $D^{-}K^{+}$ and $D^{0}K^{0}_{S}$, and the latter two are nearly equal indeed. This behavior might be attributed to the difference between rapidity distributions of $D^{+}$ and $D^{-}$ in the FHS. Meanwhile, the yield of $T_{cs0}^{*}(2870)^{0}$ increases with the increasing of the $R_{0}$ range from 1.0 fm $<R_{0}<$ 2.0 to 1.0 fm $<R_{0}<$ 5.0 fm, and to 1.0 fm $<R_{0}<$ 10.0 fm, as expected.
Figure 2: The simulated $y$ (left panel) and $p_{T}$ (right panel) distributions of three $T_{cs0}^{*}(2870)^{0}$ candidates recombined from the $D^{0}K^{0}_{S}$ (triangle-downs), the $D^{+}K^{-}$ (triangle-ups), and the $D^{-}K^{+}$ (circles) in $pp$ collisions at $\sqrt{s}$ = 7 TeV are shown, respectively. The black, blue, and red markers referred to results calculated in $R_{0}$ range of 1.0 fm $<R_{0}<$ 10.0 fm, 1.0 fm $<R_{0}<$ 5.0 fm, and 1.0 fm $<R_{0}<$ 2.0 fm, respectively.
Figure 3: The angular distribution between two component mesons (left panel, where triangle-ups for ($D^{+}$, $K^{-}$), circles for ($D^{-}$, $K^{+}$), triangle-downs for ($D^{0}$, $K^{0}_{s}$)), as well as the one between $D$ component meson and $T_{cs0}^{*}(2870)^{0}$ (right panel, where triangle-ups for ($D^{+}$, $T_{cs0}^{*}(2870)^{0}$), circles for ($D^{-}$, $T_{cs0}^{*}(2870)^{0}$), triangle-downs for ($D^{0}$, $T_{cs0}^{*}(2870)^{0}$)) in $pp$ collisions at $\sqrt{s}$=7 TeV. The black, blue, and red markers refer to the results calculated in the $R_{0}$ range of 1.0 fm $<R_{0}<$ 10.0 fm, 1.0 fm $<R_{0}<$ 5.0 fm, and 1.0 fm $<R_{0}<$ 2.0 fm, respectively.
Meanwhile, we calculate the angular distribution between two component mesons
$\cos\theta=\frac{\vec{p}_{1}.\vec{p}_{2}}{|\vec{p}_{1}||\vec{p}_{2}|},$ (5)
where $\vec{p}_{1}$ ($\vec{p}_{2}$) denotes the three momentum of $D$ ($K$) component meson. Similarly, the angular distribution between $D$ component meson and $T_{cs0}^{*}(2870)^{0}$ can be calculated. The above two angular distributions are given in Fig. 3. In this figure, the black, blue, and red markers are the results calculated in the $<R_{0}<$ range of 1.0 $<R_{0}<$ 10.0 fm, 1.0 $<R_{0}<$ 5.0 fm, and 1.0 $<R_{0}<$ 2.0 fm, respectively. The triangle-up, the circle, and the triangle-down refer to, respectively, the results of the angle between $D^{+}$ and $K^{-}$, between $D^{-}$ and $K^{+}$, as well as between $D^{0}$ and $K^{0}_{S}$ in the left panel. In the right panel, they refer to the results of the angle between $D^{+}$ and $T_{cs0}^{*}(2870)^{0}$, $D^{-}$ and $T_{cs0}^{*}(2870)^{0}$, as well as $D^{0}$ and $T_{cs0}^{*}(2870)^{0}$, respectively.
In the left panel of Fig. 3 we observe the yield of $T_{cs0}^{*}(2870)^{0}$ is the largest (smallest) if the angle between momentum vectors of $D$ and $K$ component meson pair is close to $0^{\circ}$ ($\cos\theta$ = 1) (if the angle between momentum vectors of $D$ and $K$ component meson pair is close to $180^{\circ}$ ($\cos\theta$ = -1)). It indicates the $T_{cs0}^{*}(2870)^{0}$ is easier to produce if the momentum vectors of $D$ and $K$ meson pair are convergent, as expected. The similar behavior is also observed in the right panel. The Fig. 3 shows also that the yield of $T_{cs0}^{*}(2870)^{0}$ increases with increasing $R_{0}$ range from 1.0 fm $<R_{0}<$ 2.0 fm, to 1.0 fm $<R_{0}<$ 5.0 fm, and to 1.0 fm $<R_{0}<$ 10.0 fm.
4 Summary
In this letter, we successfully recombined three sets of $T_{cs0}^{*}(2870)^{0}$ molecular state candidates based on the PACIAE model simulated final hadronic state in $pp$ collisions at $\sqrt{s}$ = 7 TeV. Our results support the LHCb observation of the $T_{cs0}^{*}(2870)^{0}$ molecular state in the decay mode of $B^{-}\to D^{-}D^{0}K^{0}_{s}$. We predict the $y$ distributions, the $p_{T}$ spectra, and the angular distributions for the three sets of $T_{cs0}^{*}(2870)^{0}$ molecular state candidates, which are worthy to be experimentally measured in future.
The extended studies for the $pp$ collisions at $\sqrt{s}$ = 8 and 13 TeV and for the $p-Pb$ and/or the $Pb-Pb$ collisions at LHC energies are required. Other observables, such as elliptic flow, etc. are also worthy to be investigated.
Acknowledgements
This work is supported by the National Natural Science Foundation of China under Grants No. 12375135. Y.L.Y. acknowledges the Continuous Basic Scientific Research Project (Grant No. WDJC-2019-13). The work of W.C.Z. is supported by the Natural Science Basic Research Plan in Shaanxi Province of China(Program No. 2023-JCYB-012). H.Z. acknowledges the financial support from Key Laboratory of Quark and Lepton Physics in Central China Normal University under grant No. QLPL2024P01.

### v2

@@
 Abstract
-We have used the parton and hadron cascade model PACIAE together with the Dynamically Constrained Phase-space Coalescence model (DCPC) to study the $T_{cs0}^{*}(2870)^{0}$ production in $pp$ collision at $\sqrt{s}$ = 7 TeV, in line with the LHCb observation of $T_{cs0}^{*}(2870)^{0}$ in the $B^{-}\to D^{-}D^{0}K^{0}_{S}$ decays in $pp$ collisions at $\sqrt{s}$ = 7, 8, and 13 TeV [PRL 134(2025)101901].
+We have used the parton and hadron cascade model PACIAE together with the Dynamically Constrained Phase-space Coalescence model (DCPC) to study the $T_{cs0}^{*}(2870)^{0}$ production in $pp$ collision at $\sqrt{s}$ = 7 TeV, following the LHCb observation of $T_{cs0}^{*}(2870)^{0}$ in the $B^{-}\to D^{-}D^{0}K^{0}_{S}$ decays in $pp$ collisions at $\sqrt{s}$ = 7, 8, and 13 TeV [PRL 134(2025)101901].
 The final hadronic states of the $pp$ collisions at $\sqrt{s}$ = 7 TeV are first simulated by the PACIAE model.
-Three sets of $T_{cs0}^{*}(2870)^{0}$ candidates are then recombined by the DCPC model using the constituent meson pair of $D^{0}K^{0}_{S}$, $D^{+}K^{-}$, and $D^{-}K^{+}$ based on the above simulated final hadronic states, respectively.
+Four sets of $T_{cs0}^{*}(2870)^{0}$ candidates are then recombined by the DCPC model using the constituent meson pair of $D^{0}K^{0}_{S}$, $D^{+}K^{-}$, $D^{*+}K^{*-}$, and $D^{*0}\bar{K}^{*0}$ based on the above simulated final hadronic states, respectively.
 We calculate their rapidity distributions, transverse momentum spectra, and angular distribution between the two component mesons, as well as angular distribution between $D$ component meson and $T_{cs0}^{*}(2870)^{0}$.
-Our results show that the yields of three $T_{cs0}^{*}(2870)^{0}$ candidates follow the magnitude order of $D^{+}K^{-}$ $>$ $D^{-}K^{+}$ $\sim$ $D^{0}K^{0}_{S}$ (the latter two are nearly equal indeed).
+Our results show that the yields of four $T_{cs0}^{*}(2870)^{0}$ candidates follow the magnitude order of $D^{0}K^{0}_{S}$ >> $D^{+}K^{-}$ >> $D^{*+}K^{*-}$ $\sim$ $D^{*0}\bar{K}^{*0}$.
 Similar ordering behavior is also observed in the aforementioned distributions.
@@
 […]
-[third]organization=School of Physics and Information Technology, Shaanxi Normal University,city=Xi’an, postcode=710119, country=China
+[third]organization=School of Physics and Electronic Science, Guizhou Normal University,city=Guiyang, postcode=550025, country=China
 […]
-[fourth]organization=China Institute of Atomic Energy, P. O. Box 275 (10),city=Beijing, postcode=102413, country=China
+[fourth]organization=School of Physics and Information Technology, Shaanxi Normal University,city=Xi’an, postcode=710119, country=China
+[…]
+[fifth]organization=China Institute of Atomic Energy, P. O. Box 275 (10),city=Beijing, postcode=102413, country=China
 1 Introduction
@@
 In Refs.
-[15, 16, 17, 18, 19, 20, 21, 22] they have been suggested as compact tetraquark states.
+[15, 16, 17, 18, 19, 20, 21, 22, 23] they have been suggested as compact tetraquark states.
 But they were interpreted as hadronic molecular states in Refs.
 [24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34].
-In line with Ref.
-[14], we assume $T_{cs0}^{*}(2870)^{0}$ is a hadronic molecular state and employ the parton and hadron cascade model PACIAE [33] together with the Dynamically Constrained Phase-space Coalescence model (DCPC [34]; PACIAE + DCPC) to study its nature and properties.
+The relative momentum between the hadrons in a hadronic molecule is much smaller than the hadron mass, making it suitable for analysis with non-relativistic methods  [35].
+The nature of exotic states as hadronic molecules can be well accounted for by the coalescence mechanism [36].
+Therefore, in this work, we assume $T_{cs0}^{*}(2870)^{0}$ produced in the final hadronic states (FHS) is a hadronic molecular state and employ the parton and hadron cascade model PACIAE [37] together with the Dynamically Constrained Phase-space Coalescence model (DCPC [38]; PACIAE + DCPC) to study its nature and properties.
 The PACIAE + DCPC model has been successfully used to describe the production of light nuclei (such as deuterons and hypertritons) in heavy-ion collisions at various center-of-mass energies [38, 39].
 It has also been applied to describe the production of exotic states, e.g., X(3872) [40, 41, 42], $P_{c}$ states [43], $Z_{c}^{\pm}$(3900) [44], X(2370) [45], and G(3900) [46] in nuclear collisions.
-Figure 1: A schematic diagram of PACIAE simulation for high-energy proton-proton collisions.
 In this letter, the PACIAE model is first employed to simulate the final hadronic states (FHS) in $pp$ collisions at $\sqrt{s}$ = 7 TeV.
-The $T_{cs0}^{*}(2870)^{0}$ molecular state is then coalesced in this FHS with the component meson pair of $D^{0}K^{0}_{S}$ and $D^{+}K^{-}$ (with conjugate $D^{-}K^{+}$), by the DCPC model.
+The $T_{cs0}^{*}(2870)^{0}$ molecular state is then coalesced in this FHS with the component meson pair of $D^{0}K^{0}_{S}$ and $D^{+}K^{-}$, by the DCPC model.
 These two component meson pairs correspond to the $T_{cs0}^{*}(2870)^{0}$ candidates observed by LHCb in decay channels of $B^{-}\to D^{-}D^{0}K^{0}_{S}$ and $B^{-}\to D^{-}D^{+}K^{-}$, respectively.
-The three sets of $T_{cs0}^{*}(2870)^{0}$ candidates, composed, respectively, of $D^{0}K^{0}_{S}$, $D^{+}K^{-}$, and $D^{-}K^{+}$ have the same quark content of $cs\bar{u}\bar{d}$.
-We calculate their rapidity ($y$) distributions, transverse momentum ($p_{T}$) spectra, and angular distribution between two component mesons of $D$ and $K$ ($D^{0}$ and $K^{0}_{S}$, $D^{+}$ and $K^{-}$, $D^{-}$ and $K^{+}$), as well as the angular distribution between $D$ ($D^{0}$, $D^{+}$, or $D^{-}$) component meson and $T_{cs0}^{*}(2870)^{0}$.
-Our results show that the yield of $T_{cs0}^{*}(2870)^{0}$ composed of $D^{+}K^{-}$ is larger the ones composed of $D^{-}K^{+}$ and $D^{0}K^{0}_{S}$.
-The latter two are nearly equal indeed.
+The four sets of $T_{cs0}^{*}(2870)^{0}$ candidates composed of $D^{0}K^{0}_{S}$, $D^{+}K^{-}$, $D^{*+}K^{*-}$, and $D^{*0}\bar{K}^{*0}$ have the same quark content of $cs\bar{u}\bar{d}$.
+We calculate their rapidity ($y$) distributions, transverse momentum ($p_{T}$) spectra.
+Then we compute the angular distribution between two component mesons of $D$ and $K$ ($D^{0}$ and $K^{0}_{S}$, $D^{+}$ and $K^{-}$), as well as the angular distribution between $D$ ($D^{0}$ or $D^{+}$) component meson and $T_{cs0}^{*}(2870)^{0}$.
+Our results show that the yield of $T_{cs0}^{*}(2870)^{0}$ composed of $D^{0}K^{0}_{S}$ is larger than the one composed of $D^{+}K^{-}$, while the yields of $T_{cs0}^{*}(2870)^{0}$ composed of $D^{*+}K^{*-}$ and $D^{*0}\bar{K}^{*0}$ are lower than the former two and almost equal.
 Similar ordering behavior is also found in all the aforementioned distributions.
@@
 The parton and hadron cascade model PACIAE [47, 48] based on PYTHIA [49], can be used to describe high-energy elementary particle collisions as well as nucleus-nucleus collisions.
-It contains four stages of the parton initialization, the parton rescattering, the hadronization, and the hadron rescattering, as shown in Fig.
-1 for a $pp$ collision.
+It contains four stages of the parton initialization, the partonic rescattering, the hadronization, and the hadronic rescattering.
 The latest version of PACIAE 4.0 [37] is based on PYTHIA 8.3 [50].
@@
 Then the gluons split and the energetic quarks (antiquarks) deexcitation is implemented.
-In the next stage of parton rescattering, the leading-order perturbative quantum chromodynamics (LO-pQCD) parton-parton scattering cross section [47] is employed to supply the final partonic state (FPS) composed of abundant quarks and antiquarks with their four-coordinate and four-momentum.
-The hadronization process is then achieved via the Lund string fragmentation mechanism and/or the phenomenological coalescence model [33], and the latter is chosen in this letter.
-Eventually, the final hadronic state (FHS) is produced after the hadron rescattering.
+In the next stage of partonic rescattering, the leading-order perturbative quantum chromodynamics (LO-pQCD) parton-parton scattering cross section [51] is employed to supply the final partonic state (FPS) composed of abundant quarks and antiquarks with their four-coordinate and four-momentum.
+The hadronization process is then achieved via the Lund string fragmentation mechanism and/or the phenomenological coalescence model [37], and the former is chosen in this letter.
+Eventually, the final hadronic state (FHS) is produced after the hadronic rescattering.
 This FHS is composed of numerous hadrons with their four-coordinate and four-momentum.
+In the process of hadronic rescattering, there are hundreds of inelastic reaction channels of light hadrons, and more inelastic reactions involving the heavy hadrons are included [37].
+Here the reaction channels of $D$ and $K$ mesons are listed as an example:
+$D+\pi$ $\rightleftharpoons D^{*}+\rho,$ $\quad D^{*}+\pi$ $\rightleftharpoons D+\rho,$
+$D+N$ $\rightleftharpoons D^{*}+N,$ $\quad D^{*}+\bar{N}$ $\rightleftharpoons D+\bar{N},$
+$\bar{K}N$ $\rightleftharpoons\pi Y,$ $\quad K\bar{N}$ $\rightleftharpoons\pi\bar{Y},$
+$\bar{K}Y$ $\rightleftharpoons\pi\Xi,$ $\quad K\bar{Y}$ $\rightleftharpoons\pi\bar{\Xi},$
+$\bar{K}N$ $\rightleftharpoons K\Xi,$ $\quad K\bar{N}$ $\rightleftharpoons\bar{K}\bar{\Xi},$
+$K\bar{\Xi}$ $\rightleftharpoons\pi\bar{\Omega},$ $\quad\bar{K}\Xi$ $\rightleftharpoons\pi\Omega,$
+where $Y$ stands for $\Lambda$ or $\Sigma$.
 2.2 DCPC model recombining $T_{cs0}^{*}(2870)^{0}$ candidates
 The DCPC model is based on the quantum statistical mechanics [52, 53], where the yield of $N$-particle cluster is defined as
-$Y_{1}=\int...\int_{E_{a}\leq H\leq E_{b}}\frac{\mathrm{d}\vec{q}_{1}\mathrm{d}\vec{p}_{1}...\mathrm{d}\vec{q}_{N}\mathrm{d}\vec{p}_{N}}{h^{3N}},$ (1)
+$Y_{1}=\int...\int_{E_{a}\leq H\leq E_{b}}\frac{\mathrm{d}\vec{r}_{1}\mathrm{d}\vec{p}_{1}...\mathrm{d}\vec{r}_{N}\mathrm{d}\vec{p}_{N}}{h^{3N}},$ (1)
 where $E_{a}$ and $E_{b}$ denote the lower and upper energy thresholds of the cluster.
-The $\vec{q}_{i}$ and $\vec{p}_{i}$ are, respectively, the coordinate and momentum of the $i$-th particle.
+The $\vec{r}_{i}$ and $\vec{p}_{i}$ are, respectively, the coordinate and momentum of the $i$-th particle.
 In order for the cluster to exist naturally, some dynamical constraints are introduced in the DCPC model.
 Thus the yield of $T_{cs0}^{*}(2870)^{0}$ composed of $D$ and $K$ component meson pair, for instance, is calculated by
-$Y_{T_{cs0}^{*}(2870)^{0}}=\int...\int\delta_{12}\frac{\mathrm{d}\vec{q}_{1}\mathrm{d}\vec{p}_{1}\mathrm{d}\vec{q}_{2}\mathrm{d}\vec{p}_{2}}{h^{6}},$ (2)
+$Y_{T_{cs0}^{*}(2870)^{0}}=\int...\int\delta_{12}\frac{\mathrm{d}\vec{r}_{1}\mathrm{d}\vec{p}_{1}\mathrm{d}\vec{r}_{2}\mathrm{d}\vec{p}_{2}}{h^{6}},$ (2)
 where
-$\delta_{12}=\begin{cases}1\quad\rm{if}\quad 1\equiv D,2\equiv K,\\ \qquad m_{0}-\Delta m\leq m_{\rm{inv}}\leq m_{0}+\Delta m,\\ \qquad R_{i}\leq R_{0},(i=1,2)\\ 0\quad\rm{otherwise}.\end{cases}$ (3)
+$\delta_{12}=\begin{cases}1\quad\rm{if}\quad 1\equiv D,2\equiv K,\\ \qquad m_{0}-\Delta m\leq m_{\rm{inv}}\leq m_{0}+\Delta m,\\ \qquad r_{i0}\leq R_{0},(i=1,2)\\ 0\quad\rm{otherwise}.\end{cases}$ (3)
 In Eq.(3), $m_{0}$ ($\Delta m$) denotes the mass (mass uncertainty parameter) of the cluster ($T_{cs0}^{*}(2870)^{0}$ candidate).
-$R_{0}$ and $R_{i}$ refer to the radius of the cluster (a free parameter) and the relative distance between the $i$-th component meson and the center of mass of the cluster, respectively.
-The invariant mass $m_{inv}$ is defined as
+$R_{0}$ and $r_{i0}$ refer to the radius of the cluster (a free parameter) and the relative distance between the $i$-th component meson and the center of mass of the cluster, respectively.
+The invariant mass $m_{\rm{inv}}$ is defined as
 $m_{\rm{inv}}=\sqrt{(E_{1}+E_{2})^{2}-(\vec{p}_{1}+\vec{p}_{2})^{2}},$ (4)
 where $E_{1}$ ($E_{2}$) and $\vec{p}_{1}$ ($\vec{p}_{2}$) denote, respectively, the energy and momentum of $D$ ($K$) component meson.
+Figure 1: The yield of $T_{cs0}^{*}(2870)^{0}$ recombined from $D^{0}K^{0}_{S}$ (triangles-down), $D^{+}K^{-}$ (triangles-up), $D^{*+}K^{*-}$ (circles), and $D^{*0}\bar{K}^{*0}$ (squares) varying with the parameter $\Delta m$ under the $R_{0}$ range of 1.0 fm $R_{0}<$ 5.0 fm in $pp$ collisions at $\sqrt{s}$ = 7 TeV.
 2.3 Calculation routine in DCPC
-We first construct a $T_{cs0}^{*}(2870)^{0}$ component meson list (consisted of $K^{+}$, $K^{-}$, $K^{0}_{s}$, $D^{+}$, $D^{-}$, and $D^{0}$) based on the hadron list in the FHS generated by the PACIAE model.
-According to the Eq.
-(3) we then construct a two-layer circulations (loops) over the component meson list to filter out any one of possible candidate among the studied three sets of $T_{cs0}^{*}(2870)^{0}$ candidates (composed of $D^{0}K^{0}_{S}$, $D^{+}K^{-}$, and $D^{-}K^{+}$).
-Here the parameters in Eq.
-(3) are set as follows: $m_{0}=2866$ MeV/$c^{2}$ (the experimentally observed mass peak of $T_{cs0}^{*}(2870)^{0}$ molecular state); $\Delta m$ equals to 28.5 MeV/$c^{2}$ (the half decay width of the $T_{cs0}^{*}(2870)^{0}$ [50]); and the $R_{0}$ is assumed to be in the range ($R_{0}$ range) of 1.0 fm $<R_{0}<$ 10.0 fm  [51, 52], or 1.0 fm $<R_{0}<$ 5.0 fm, or 1.0 fm $<R_{0}<$ 2.0 fm.
+[…]
 3 Results and discussion
-The PACIAE 4.0 model (with coalescence hadronization) is used to generate 300 million final hadronic states for the $pp$ collisions at $\sqrt{s}$ = 7 TeV.
-Here we adjust only two main parameters of adj1(16) (the allowed number of deexcitation generations of the energetic quark (antiquark)) and adj1(17) (the threshold energy of deexcitation).
+The PACIAE 4.0 model (with string fragmentation) is used to generate 800 million final hadronic states for the $pp$ collisions at $\sqrt{s}$ = 7 TeV.
+Here we adjust only two main parameters of adj1(10) (the K factor multiplying the differential cross sections for hard parton-parton process) and adj1(34) (the width of $p_{T}$ sampling).
 Other parameters are default.
-The adj1(16) and adj1(17) are fixed by fitting the ALICE measured $\pi^{+}$, $K^{+}$, $D^{0}$ and $D^{+}$ yields at mid-rapidity region in $pp$ collisions at $\sqrt{s}$ = 7 TeV [53, 54].
-The simulated yields (with fitted parameters of adj1(16)=1 and adj1(17)=2.15 GeV) and the corresponding experimental data are shown in Table 1.
-Then we use above generated final hadronic states to recompose three sets of $T_{cs0}^{*}(2870)^{0}$ hadronic molecular state candidates by the DCPC model using the constituent meson pairs of $D^{0}K^{0}_{S}$, $D^{+}K^{-}$, and $D^{-}K^{+}$, respectively.
-Table 1: The yields of $\pi^{+}$, $K^{+}$, $D^{0}$, and $D^{+}$ simulated by PACIAE model, compared with the ALICE data at mid-rapidity region in $pp$ collisions at $\sqrt{s}$=7 TeV [53, 54].
+The adj1(10) and adj1(34) are fixed by fitting the ALICE measured $p_{T}$ spectra of $\pi^{+}+\pi^{-}$, $K^{+}+K^{-}$, $K^{0}_{S}$, $(D^{0}+\bar{D}^{0})/2$ and $(D^{+}+D^{-})/2$ at mid-rapidity region in $pp$ collisions at $\sqrt{s}$ = 7 TeV [56, 57, 58].
+The simulated $p_{T}$ spectra (with adj1(10)=0.8 and adj1(34)=0.46) and the corresponding experimental data are shown in Fig 2.
+The simulated yields of $\pi^{+}$, $K^{-}$, $K^{0}_{S}$, $D^{0}$ and $D^{+}$ and the corresponding experimental data are also presented in Table 1.
+Table 1: The yields of $\pi^{+}$, $K^{-}$, $K^{0}_{S}$, $D^{0}$ and $D^{+}$ simulated by PACIAE model, compared with the ALICE data at mid-rapidity region in $pp$ collisions at $\sqrt{s}$=7 TeV [56, 57, 58].
+Here the yields of $D^{0}$ and $D^{+}$ are prompt production (average of particles and antiparticles).
 Particle ALICE PACIAE
-$\pi^{+}$ $2.26\pm 0.10$ 2.27
-$K^{+}$ $0.286\pm 0.016$ 0.284
-$D^{0}$ $(8.04\pm 1.57)\times 10^{-3}$ $8.25\times 10^{-3}$
-$D^{+}$ $(2.93\pm 0.72)\times 10^{-3}$ $3.06\times 10^{-3}$
-The simulated $y$ (left panel) and $p_{T}$ (right panel) distributions of three $T_{cs0}^{*}(2870)^{0}$ candidates (recombined by $D^{0}K^{0}_{S}$, $D^{+}K^{-}$, and $D^{-}K^{+}$) in $pp$ collisions at $\sqrt{s}$ = 7 TeV are given in Fig.
-2.
-In this figure, the black, blue, and red markers are the results calculated in the $R_{0}$ range of 1.0 $<R_{0}<$ 10.0 fm, 1.0 $<R_{0}<$ 5.0 fm, and 1.0 $<R_{0}<$ 2.0 fm, respectively.
-The triangle-up, the circle, and the triangle-down refer to, respectively, the results of the $D^{+}K^{-}$, the $D^{-}K^{+}$, and the $D^{0}K^{0}_{s}$ component meson pairs.
-It is observed that the yield of $T_{cs0}^{*}(2870)^{0}$ composed of $D^{+}K^{-}$ is larger than the ones composed of $D^{-}K^{+}$ and $D^{0}K^{0}_{S}$, and the latter two are nearly equal indeed.
-This behavior might be attributed to the difference between rapidity distributions of $D^{+}$ and $D^{-}$ in the FHS.
-Meanwhile, the yield of $T_{cs0}^{*}(2870)^{0}$ increases with the increasing of the $R_{0}$ range from 1.0 fm $<R_{0}<$ 2.0 to 1.0 fm $<R_{0}<$ 5.0 fm, and to 1.0 fm $<R_{0}<$ 10.0 fm, as expected.
-Figure 2: The simulated $y$ (left panel) and $p_{T}$ (right panel) distributions of three $T_{cs0}^{*}(2870)^{0}$ candidates recombined from the $D^{0}K^{0}_{S}$ (triangle-downs), the $D^{+}K^{-}$ (triangle-ups), and the $D^{-}K^{+}$ (circles) in $pp$ collisions at $\sqrt{s}$ = 7 TeV are shown, respectively.
-The black, blue, and red markers referred to results calculated in $R_{0}$ range of 1.0 fm $<R_{0}<$ 10.0 fm, 1.0 fm $<R_{0}<$ 5.0 fm, and 1.0 fm $<R_{0}<$ 2.0 fm, respectively.
-Figure 3: The angular distribution between two component mesons (left panel, where triangle-ups for ($D^{+}$, $K^{-}$), circles for ($D^{-}$, $K^{+}$), triangle-downs for ($D^{0}$, $K^{0}_{s}$)), as well as the one between $D$ component meson and $T_{cs0}^{*}(2870)^{0}$ (right panel, where triangle-ups for ($D^{+}$, $T_{cs0}^{*}(2870)^{0}$), circles for ($D^{-}$, $T_{cs0}^{*}(2870)^{0}$), triangle-downs for ($D^{0}$, $T_{cs0}^{*}(2870)^{0}$)) in $pp$ collisions at $\sqrt{s}$=7 TeV.
-The black, blue, and red markers refer to the results calculated in the $R_{0}$ range of 1.0 fm $<R_{0}<$ 10.0 fm, 1.0 fm $<R_{0}<$ 5.0 fm, and 1.0 fm $<R_{0}<$ 2.0 fm, respectively.
-Meanwhile, we calculate the angular distribution between two component mesons
+$\pi^{+}$ $2.26\pm 0.10$ 2.17
+$K^{-}$ $0.286\pm 0.016$ 0.297
+$K^{0}_{S}$ $0.280\pm 0.015$ 0.300
+$D^{0}$ $(6.83\pm 0.77)\times 10^{-3}$ $6.38\times 10^{-3}$
+$D^{+}$ $(3.10\pm 0.84)\times 10^{-3}$ $2.99\times 10^{-3}$
+Then we use above generated final hadronic states to recompose four sets of $T_{cs0}^{*}(2870)^{0}$ hadronic molecular state candidates by the DCPC model using the constituent meson pairs of $D^{0}K^{0}_{S}$, $D^{+}K^{-}$, $D^{*+}K^{*-}$, and $D^{*0}\bar{K}^{*0}$, respectively.
+Figure 2: The simulated $p_{T}$ spectra and the corresponding experimental data of $\pi^{+}+\pi^{-}$, $K^{+}+K^{-}$, $K^{0}_{S}$, $(D^{0}+\bar{D}^{0})/2$ and $(D^{+}+D^{-})/2$.
+The empty markers refer to the results simulated by the PACIAE model, and the solid markers refer to the ALICE data [56, 57, 58].
+The simulated $y$ (left panel) and $p_{T}$ (right panel) distributions of four $T_{cs0}^{*}(2870)^{0}$ candidates (recombined by $D^{0}K^{0}_{S}$, $D^{+}K^{-}$, $D^{*+}K^{*-}$, and $D^{*0}\bar{K}^{*0}$) in $pp$ collisions at $\sqrt{s}$ = 7 TeV are given in Fig.
+3.
+In this figure, the black, blue, and red markers are the results calculated in the $R_{0}$ range of 1.0 $R_{0}<$ 5.0 fm, 1.0 $R_{0}<$ 3.0 fm, and 1.0 $R_{0}<$ 2.0 fm, respectively.
+The triangles-down, the triangles-up, the circles, and the squares refer to, respectively, the results of the $D^{0}K^{0}_{S}$, $D^{+}K^{-}$, $D^{*+}K^{*-}$, and $D^{*0}\bar{K}^{*0}$ component meson pairs.
+It is observed that the yield of $T_{cs0}^{*}(2870)^{0}$ composed of $D^{0}K^{0}_{S}$ is larger than the one composed of $D^{+}K^{-}$, which is consistent with the yields of the constituent particles of $T_{cs0}^{*}(2870)^{0}$ in Table 1.
+The yields of $T_{cs0}^{*}(2870)^{0}$ composed of $D^{*+}K^{*-}$ and $D^{*0}\bar{K}^{*0}$ are nearly equal, which is attributed to the similar yields of their constituent mesons $D^{*+}$ and $D^{*0}$ as reported in Ref.
+[59].
+The yields of $T_{cs0}^{*}(2870)^{0}$ composed of $D^{*+}K^{*-}$ and $D^{*0}\bar{K}^{*0}$ are lower than the former two.
+It can be understood by the fact that the productions of inclusive $D^{+}$ ($D^{0}$) and $K^{-}$ ($K^{0}_{S}$) are larger than those of $D^{*+}$ ($D^{*0}$) and $K^{*-}$ ($\bar{K}^{*0}$).
+Meanwhile, the yield of $T_{cs0}^{*}(2870)^{0}$ increases with the increasing of the $R_{0}$ range from 1.0 fm $R_{0}<$ 2.0 to 1.0 fm $R_{0}<$ 3.0 fm, and to 1.0 fm $R_{0}<$ 5.0 fm, as expected.
+Figure 3: The simulated $y$ (left panel) and $p_{T}$ (right panel) distributions of four $T_{cs0}^{*}(2870)^{0}$ candidates recombined from the $D^{0}K^{0}_{S}$ (triangles-down), the $D^{+}K^{-}$ (triangles-up), the $D^{*+}K^{*-}$ (circles), and the $D^{*0}\bar{K}^{*0}$ (squares) in $pp$ collisions at $\sqrt{s}$ = 7 TeV are shown, respectively.
+The black, blue, and red markers refer to results calculated in $R_{0}$ range of 1.0 fm $R_{0}<$ 5.0 fm, 1.0 fm $R_{0}<$ 3.0 fm, and 1.0 fm $R_{0}<$ 2.0 fm, respectively.
+Figure 4: The angular distribution between two component mesons (left panel, where triangles-down for ($D^{0}$, $K^{0}_{s}$) and triangles-up for ($D^{+}$, $K^{-}$)), as well as the one between $D$ component meson and $T_{cs0}^{*}(2870)^{0}$ (right panel, where triangles-down for ($D^{0}$, $T_{cs0}^{*}(2870)^{0}$) and triangles-up for ($D^{+}$, $T_{cs0}^{*}(2870)^{0}$)) in $pp$ collisions at $\sqrt{s}$=7 TeV.
+The black, blue, and red markers refer to the results calculated in the $R_{0}$ range of 1.0 fm $R_{0}<$ 5.0 fm, 1.0 fm $R_{0}<$ 3.0 fm, and 1.0 fm $R_{0}<$ 2.0 fm, respectively.
+We also calculate the angular distribution between two component mesons
 $\cos\theta=\frac{\vec{p}_{1}.\vec{p}_{2}}{|\vec{p}_{1}||\vec{p}_{2}|},$ (5)
-where $\vec{p}_{1}$ ($\vec{p}_{2}$) denotes the three momentum of $D$ ($K$) component meson.
+where $\vec{p}_{1}$ ($\vec{p}_{2}$) denotes the momentum of $D$ ($K$) component meson.
 Similarly, the angular distribution between $D$ component meson and $T_{cs0}^{*}(2870)^{0}$ can be calculated.
 The above two angular distributions are given in Fig.
-3.
-In this figure, the black, blue, and red markers are the results calculated in the $<R_{0}<$ range of 1.0 $<R_{0}<$ 10.0 fm, 1.0 $<R_{0}<$ 5.0 fm, and 1.0 $<R_{0}<$ 2.0 fm, respectively.
-The triangle-up, the circle, and the triangle-down refer to, respectively, the results of the angle between $D^{+}$ and $K^{-}$, between $D^{-}$ and $K^{+}$, as well as between $D^{0}$ and $K^{0}_{S}$ in the left panel.
-In the right panel, they refer to the results of the angle between $D^{+}$ and $T_{cs0}^{*}(2870)^{0}$, $D^{-}$ and $T_{cs0}^{*}(2870)^{0}$, as well as $D^{0}$ and $T_{cs0}^{*}(2870)^{0}$, respectively.
+4.
+Note that the results of the angular distribution between $D^{*+}$ and $K^{*-}$, $D^{*0}$ and $\bar{K}^{*0}$, as well as $D^{*+}$ ($D^{*0}$) and $T_{cs0}^{*}(2870)^{0}$ are not shown here due to their immense computational demands.
+In this figure, the black, blue, and red markers refer to the results calculated in the $R_{0}$ range of 1.0 $R_{0}<$ 5.0 fm, 1.0 $R_{0}<$ 3.0 fm, and 1.0 $R_{0}<$ 2.0 fm, respectively.
+The triangles-down and the triangles-up refer to, respectively, the results of the angle between $D^{0}$ and $K^{0}_{S}$, as well as between $D^{+}$ and $K^{-}$ in the left panel.
+In the right panel, they refer to the results of the angle between $D^{0}$ and $T_{cs0}^{*}(2870)^{0}$, as well as $D^{+}$ and $T_{cs0}^{*}(2870)^{0}$, respectively.
 In the left panel of Fig.
@@
 The Fig.
-3 shows also that the yield of $T_{cs0}^{*}(2870)^{0}$ increases with increasing $R_{0}$ range from 1.0 fm $<R_{0}<$ 2.0 fm, to 1.0 fm $<R_{0}<$ 5.0 fm, and to 1.0 fm $<R_{0}<$ 10.0 fm.
+4 shows also that the yield of $T_{cs0}^{*}(2870)^{0}$ increases with increasing $R_{0}$ range from 1.0 fm $R_{0}<$ 2.0 fm, to 1.0 fm $R_{0}<$ 3.0 fm, and to 1.0 fm $R_{0}<$ 5.0 fm.
 4 Summary
-In this letter, we successfully recombined three sets of $T_{cs0}^{*}(2870)^{0}$ molecular state candidates based on the PACIAE model simulated final hadronic state in $pp$ collisions at $\sqrt{s}$ = 7 TeV.
-Our results support the LHCb observation of the $T_{cs0}^{*}(2870)^{0}$ molecular state in the decay mode of $B^{-}\to D^{-}D^{0}K^{0}_{s}$.
-We predict the $y$ distributions, the $p_{T}$ spectra, and the angular distributions for the three sets of $T_{cs0}^{*}(2870)^{0}$ molecular state candidates, which are worthy to be experimentally measured in future.
+In this letter, we successfully recombined four sets of $T_{cs0}^{*}(2870)^{0}$ hadronic molecular state candidates based on the PACIAE model simulated final hadronic state in $pp$ collisions at $\sqrt{s}$ = 7 TeV.
+We find that the yields of four $T_{cs0}^{*}(2870)^{0}$ candidates follow the magnitude order of $D^{0}K^{0}_{S}$ >> $D^{+}K^{-}$ >> $D^{*+}K^{*-}$ $\sim$ $D^{*0}\bar{K}^{*0}$.
+We predict the $y$ distributions, the $p_{T}$ spectra, and the angular distributions for the four sets of $T_{cs0}^{*}(2870)^{0}$ hadronic molecular state candidates, which are worthy to be experimentally measured in future.
 The extended studies for the $pp$ collisions at $\sqrt{s}$ = 8 and 13 TeV and for the $p$-Pb and/or the Pb-Pb collisions at LHC energies are required.